OpenStudy (anonymous):

Administrators at MG hospital believe that the hospital's expenditures E(b) measures in dollars, are a function of how many beds (B) are in use with; E(b)=1400 + (B+1)^2. On the other hand, the number of beds is a function of time t, measured in days and it is estimated that: B(t)= 20sin(t/10) + 50. At what rate are the expenditures decreasing when t =100?

ganeshie8 (ganeshie8):

use chain rule

ganeshie8 (ganeshie8):

\(\dfrac{dE}{dt} = \dfrac{dE}{dB} \times \dfrac{dB}{dt}\)

OpenStudy (anonymous):

what would dE/dB be?

ganeshie8 (ganeshie8):

\(E=1400 + (B+1)^2 \) differentiate with respect to B

OpenStudy (anonymous):

so 2(B+1) * 2cos(t/10) ?

ganeshie8 (ganeshie8):

looks good, plugin B value[B= 20sin(t/10) + 50] and evaluate at t = 100

OpenStudy (anonymous):

I'm getting -135 when I should be getting a +135

ganeshie8 (ganeshie8):

\(\dfrac{dE}{dt} = \dfrac{dE}{dB} \times \dfrac{dB}{dt} \) \(~~~~~ = 2(B+1)\times 2\cos(\frac{t}{10})\) \(~~~~~ = 2(20\sin(t/10) + 50 +1)\times 2\cos(\frac{t}{10})\)

ganeshie8 (ganeshie8):

\(\dfrac{dE}{dt}= 2(20\sin(t/10) + 50 +1)\times 2\cos(\frac{t}{10})\) \(\dfrac{dE}{dt}\Bigg |_{t=100 } = -135\)

OpenStudy (anonymous):

OH cause its decreasing so thats why its negative right?

ganeshie8 (ganeshie8):

that means expenditures are DECREASING at a rate of 135

ganeshie8 (ganeshie8):

you got it !

ganeshie8 (ganeshie8):

On number line : "-135" is same as saying "left +135 units"

OpenStudy (anonymous):

ok thanks!

ganeshie8 (ganeshie8):

np :)